OpenStudy (anonymous):

having doubt plz help lim x->0 [sin(abs(x))/x] where abs is mod and [] denotes greatest function integer for x-->0+ am getting 0 but for x-->0- how come it is -1 i know Limit won't exist getting it from 2nd method plz tell me m confused -_- see limx->0- [-sinx/x] inside part should be lesser than 1 nah :O how come the value inside would be greater than 1 :O if it greater than 1 then m getting answer but how it will be greater than 1 :'( limx->0- [-sinx/x] inside part sinx/x will be lesser than or greater than one ??????

OpenStudy (anonymous):

Now let's right it as a piece-wise function:\[\bf \frac{\sin|x|}{x}=\frac{\sin(x)}{x} \ for \ x \ge0 \ and \ \frac{\sin(-x)}{x}=-\frac{\sin(x)}{x} \ for \ x < 0\]Evaluate the limit from the right and left sides:\[\bf \lim_{\ \ x \rightarrow 0^+}\frac{\sin|x|}{x}=\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1\]\[\bf \lim_{\ \ x \rightarrow 0^-}\frac{\sin|x|}{x}=\lim_{x \rightarrow 0}-\frac{\sin(x)}{x}=-1\]Clearly, the one-sided limits are unequal hence:\[\bf \lim_{x \rightarrow 0}\frac{\sin|x|}{x}=D.N.E\] @kittycat01

OpenStudy (anonymous):

huh ? u have to take lhl and rhl nah rhl m getting but having problem with lhl

OpenStudy (anonymous):

wait that's u forgot to add greatest integer function

OpenStudy (anonymous):

What are you talking about? @kittycat01

OpenStudy (anonymous):

@genius12 -sinmod(x)/x is inside the greatest funtion interger bracket

OpenStudy (anonymous):

grrrr plz help m still having doubt thet floor bracket is confusing me :'(

OpenStudy (anonymous):

oh thats greatest integer function lol i didnt notice that

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

who can help me :'(

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

@genius12 -do u know the answer ? if no then no problem ,i m going to resolve it :)

OpenStudy (anonymous):

https://www.desmos.com/calculator/4kbaskvfsp Sorry I can't explain this, I have to leave. @kittycat01

OpenStudy (anonymous):

@genius12 -okay :)

OpenStudy (anonymous):

observe the one-sided limits are different which you can find out without graphing the function, so the limit in this case also, does not exist. @kittycat01

OpenStudy (anonymous):

@genius12 well i know limit won't exist but having difficulty in finding the value at lhl :)

OpenStudy (shubhamsrg):

for x->0 , sinx/x tends to 1-

OpenStudy (anonymous):

but it's [sinx/x] nah :P

OpenStudy (shubhamsrg):

[1-] = 0 negative of this inside [ ] will make it -1 by the same logic

OpenStudy (anonymous):

:/ []=GIF inside value is tending toward 1- lt x-0- [-sin/x] =-2 :/

OpenStudy (anonymous):

if m right then lt n->0- [n] = n-1

OpenStudy (anonymous):

*sorry n->a-

OpenStudy (shubhamsrg):

sin x /x = 0.999999 at x->0 [0.999999] = 0 [-0.99999]= -1

OpenStudy (anonymous):

wait :OOOOOOOOOOOOO

OpenStudy (anonymous):

let me think let us take this one as a example lt x->1- [-sinx/x] -sinx/x = -0.854 so inside part would be tending towards 0- so we will get -1 as a answer :/ now let us follow this for 0- lt x->0- sin(-0.1)/-0.1 =0.99 inner part is tending toward 0- so answer would be -1 -_-

OpenStudy (anonymous):

oh :OOOOO i think now m getting it

OpenStudy (anonymous):

grrr need more practise solving this type of questions -_-

OpenStudy (shubhamsrg):

hmm, practice will surely help :) joshi sir se argument kiya karo, it'll help ^^

OpenStudy (anonymous):

lol joshi sir toh pata nahi kaha hotai hai :P dikhtai hi nahi hai